OpenStudy (anonymous):

Evaluate -4x^2-2y^3 when x =1 and y = -5 FOR MEDAL PLZ HELP

OpenStudy (anonymous):

-4(1)^2-2(-5)^3 1^2 = 1 (-5)^3 = -125 -4(1)-2(-125) = -4-(-250) = 246

OpenStudy (anonymous):

thank you boss.